OpenStudy (anonymous):

Bill has to figure out the area of his square yard. He knows that one side of his yard is equal to the length of his garage. The dimensions of his rectangular garage are 19 by 10. What is the area of the yard?

OpenStudy (anonymous):

Here Right. For a square, w = h = s A = s * s The square yard has a side of 19. A = s * s = 19 * 19

OpenStudy (anonymous):

Hope this helps

OpenStudy (anonymous):

361?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks